Use the following values for mass and charge: an electron has mass me = 9.11×10-31 kg and charge -e, a proton has mass mp = 1.67×10-27 kg and charge +e, an alpha particle has mass malpha = 6.65×10-27 kg and charge +2e, where e = 1.60×10-19 C. An electron is released from rest in a vacuum between two flat, parallel metal plates that are 17.0 cm apart and are maintained at a constant electric potential difference of 730 Volts. If the electron is released at the negative plate, what is its speed just before it strikes the positive plate? What is the kinetic energy of the electron in units of eV?
Answer #1

Force, F = ma = qE

Acceleration, a = qE/m = (730/0.17)(1.6 x 10^-19)/(9.1 x 10^-31) = 7.55 x 10^14 m/s^2

Using 3rd equation of motion,

V^2 = u^2 + 2as

V^2 = 2 x 7.55 x 10^14 x 0.17

V = 1.60 x 10^7 m/s

Kinetic energy, KE = 0.5 mv^2 = 0.5 x (9.1 x 10^-31)(1.6 x 10^7)^2 = 1.17 x 10^-16 J = 730 eV

Comment in case any doubt please rate my answer.....